# Expert Router: Orchestrating Efficient Language Model Inference through Prompt Classification

### v1

@@ VI-B System-Centric Metrics
In the following, we assess the impact of increasing parallel user requests on the latency measured by the p99 response time and the system throughput. The p99 response time for the base and continuation prompt are illustrated in Figure 7 while the results of the session throughput are shown in Figure 8. In both plots, a performance improvement for the models orchestrated via the Expert Router (B-D) is visible within a range of 200 to 900 Users. To explain this improvement, it is necessary to visualize and investigate the throughput trends over time shown in Figure 9. In these plots, the throughput is calculated using the time bucket method with time windows of 60 s.
The time bucket throughput across all models is similar at low user numbers (1-100 users) as visible in Figure 9. However, as user numbers increase, the throughput of the tensor-parallelized model (A) exhibits a peak followed by a sharp decline, whereas the orchestrated models (B-D) show multiple peaks, indicating fluctuations between 200 and 2,000 tokens per second. This pattern suggests that the orchestrated models, despite the fluctuations, achieve a more stable and consistent throughput over extended periods, especially as the user count rises.
When the system receives new requests, they are handled by the batcher, which schedules them for processing. This scheduling phase can delay the computation process for generating new output tokens. This initial delay phase, visible for all systems, can be seen as a ramp-up period in Figure 9. It leads to what is referred to as ”generation stalls” a decrease in throughput observed during the initial processing stages [41]. For the tensor-parallelized model (A), peak throughput values indicate moments of optimal equilibrium between inter-GPU communication demands and batch processing. On the other hand, the orchestrated models (B-D) process multiple smaller batches concurrently over an extended period, leading to a higher average throughput and more consistent performance. This evidence supports the expert router’s effectiveness, showcasing its ability to enhance system throughput within specific operational ranges.
@@ VII Conclusion and Future Work
We introduced Expert Router, a system that efficiently orchestrates multiple specialized LLMs. Through our experimental analysis, we have demonstrated that our orchestration technique maintains throughput comparable to that of a tensor parallelized model and enhances the average system throughput in certain settings. Notably, the parameter count of the individual experts does not negatively affect the architecture’s performance. As a result, deploying larger quantized models within this framework effectively increases the total parameter count without compromising on system throughput. This advancement opens a new avenue for deploying complex model ensembles in resource-constrained environments, promising improvements in both efficiency and effectiveness.
As part of future work, we plan to create a fully unsupervised pipeline for expert creation and inference. One main focus of this approach will be to conduct a detailed analysis of the model’s ability to generate language, focusing on how different training datasets affect the performance of both the classifier and the expert models. While we used a uniform prompt distribution in this work, one could horizontally scale the number of deployed models for each cluster based on incoming prompts. This would help in keeping an even load across the used infrastructure. Another avenue for future exploration is to study if different setups for the Routing Gateway can further increase the performance. Such experiments include testing different embedding and classification algorithms or expanding clustering criteria beyond the prompt domain to incorporate user and request categories.
@@ Acknowledgment
The authors generated parts of this text with OpenAI’s language-generation models. Upon generation, the authors reviewed, edited, and revised the language.

### v2

@@ VI-B System-Centric Metrics
In the following, we assess the impact of increasing parallel user requests on the latency measured by the p99 response time and the system throughput. The suggested resilience of the Expert Router configurations to non-balanced workloads is also visible in the p99 response time trajectories shown in Figure 6 II). Both model configurations (D, E) only show minor differences between the normal and uniform request distributions. Baseline model C has the lowest p99 response time for high levels of user concurrency, caused by the previously shown stable TTFT.
To further assess the impact of Expert Router on system metrics, we present the system throughput shown in Figure 7. The size of one window in which the number of received tokens is counted is set to two seconds. For 200 and 600 concurrent users, Expert Router configuration E exhibits an almost identical throughput trajectory to baseline model C. In these scenarios, the Expert Router configuration does not negatively impact system throughput. This is further supported by the average throughput, indicated by the colored arrows on the right y-axis. It shows that both configurations achieve nearly the same average system throughput for these user counts. However, at 1,000 concurrent users, model C outperforms Expert Router configuration E. This highlights the effect of the larger TTFT values and the benefits of tensor parallelization on system throughput.
The effect of using the high parameter 70B models in the Expert Router configuration is especially visible in these plots. Across all user numbers, model D has the lowest peak throughput. Although the average throughput increases with higher user concurrency, it does not pass the 2,000 tokens/s mark. These results highlight the trade-offs involved in deploying different model configurations. When the capabilities of multiple high-parameter models, like the 70B configuration, are required, user and system metrics are more significantly affected. The less parameter-intensive configuration can achieve a similar throughput to a tensor-parallelized model. It generally does not negatively impact user and system metrics across a wide range of concurrency levels.
@@ VII Conclusion and Future Work
In this paper, we have characterized the performance of the Expert Router, a scalable routing architecture for LLM inference. The results demonstrate that the Expert Router effectively combines specialized models while maintaining robust performance in terms of throughput and latency, even with high concurrency levels and workload imbalances. Our characterization further shows that the latency introduced by the routing process remains constant even under high user concurrency, allowing model configuration, particularly the choice and tuning of parameters, to drive performance outcomes more significantly. For moderate levels of user concurrency, the inference metrics are not significantly impacted by the model size. In such scenarios, quantized high parameter models can efficiently increase the total parameter count without substantially compromising performance. At higher levels of user concurrency, smaller models can achieve very similar performance to tensor-parallelized baseline models. Such configurations require less advanced infrastructure and are especially well suited in computing environments without high-bandwidth inter-GPU connections. Conclusively, Expert Router provides a versatile solution for scalable LLM inference, adapting well to various compute environments and request scenarios while offering robust performance regardless of infrastructure limitations.
As part of future work, we plan to create a fully unsupervised pipeline for expert creation and inference. One main focus of this approach will be to conduct a detailed analysis of the model’s ability to generate language, focusing on how different training datasets affect the performance of both the classifier and the expert models. Furthermore, we plan to extend this to dynamic deployment of LoRA weights, allowing us to increase the number of experts in the systems without increasing the number of models. Another avenue for future exploration is to study if different setups for the routing gateway can further increase the performance. Such experiments include testing different embedding and classification algorithms or expanding clustering criteria beyond the prompt domain to incorporate user and task categories.
@@ Acknowledgment
The authors generated parts of this text using OpenAI’s language-generation models. Upon generation, the authors reviewed, edited, and revised the language.